OpenStudy (anonymous):

the length of the rectangle is 7 in more than double the width

OpenStudy (anonymous):

l = 7 + 2w

OpenStudy (anonymous):

ince there are two widths and two lengths, we have to double the unknowns and double the 7 inches. We want the unknowns on side side of the equal sign and the knowns on the other side. Remember that whatever we do to one side we must also do to the other side of the equal sign. Move the 14 inches (7 inches doubled) to the right side by subtracting 14 inches from both sides. Then add the 2x + 6x (which is the unknown length and unknown width doubled, because there are two widths and two lengths.) What has been done to the x? It has been multiplied by 8. To isolate the x we do the opposite of multiplication, which is division. Divide both sides by 8. That gives us x = 10 2x + 14 +6x = 94 ___- 14___= -14 8x = 80 x = 10 Let's test this. (10 + 7) + (10 + 7) + 30 + 30 = 94

OpenStudy (anonymous):

thank you for your answer but it is saying it is not right.. but thank you any way

OpenStudy (anonymous):

i meant (since)

OpenStudy (anonymous):

It is saying let w represent the w of the rectangle

OpenStudy (anonymous):

What do we know? The perimeter of a rectangle is 2 * (L+W) If we let W = width, yet unknown, then the length is 3*W+7 (from the problem statement) Perimeter of rectangle is 2* (L+W) = 2* [(3W+7) + W] Substitute in actual value for perimeter of rectangle = 94 94 = 2* [(3W+7) + W] simplify the right side: 94 = 2[4W+7] 94 = 8W+14 solve for W; compute L from W as above.

OpenStudy (anonymous):

\[The first angel of the \triangle is 17 degrees more than the second \angle. The third \angle \angle of the \triangle is three time more the second \angle. use x \to represent the unknown value. find the first \angle the second and the third\]

OpenStudy (anonymous):

Thank you great job